Joe's present age is three times Greg's age. Five years ago, the sume of their ages was 14. What is each boy's present age?
Answer by Maths68(1474)

Let
Present age of Joe = j
Present age of Greg = g


Joe's present age is three times Greg's age
j=3g..........(1)
Five years ago
Joe was = j-5
Greg was = g-5
then the sume of their ages was 14.
(j-5)+(g-5)=14
j-5+g-5=14
j+g-10=14
j+g=14+10
j+g=24..............(2)
Put the value of j from (1) to (2)
3g+g=24
4g=24
4g/4=24/4
g=6
Put the value of g in (1)
j=3g
j=3*6
j=18
Present age of Joe = j = 18 years old
Present age of Greg = g = 6 years old